# On the Use of Agentic Coding: An Empirical Study of Pull Requests on GitHub

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Integrated and visual development environmentsccs: Software and its engineering Automatic programmingccs: Software and its engineering Software evolutionccs: Software and its engineering Maintaining software
@@ 1. Introduction
Prior research has examined the impact of integrating LLM-based tools into software development workflows. For instance, Tufano et al. (Tufano et al., 2024) manually analyzed 527 commits, 327 pull requests (PRs), and 647 issues, revealing that developers employ ChatGPT for a wide range of software engineering tasks, spanning 45 distinct categories, including automating the creation or enhancement of features. Similarly, Watanabe et al. (Watanabe et al., 2024) studied the use of ChatGPT in PR reviews and the corresponding developer responses. They found that 30.7% of AI-generated suggestions during code reviews were met with skepticism or disagreement, and most of them were about the generated code.
@@
  • $RQ_{3}$: What proportion of Agentic-PRs are accepted without revisions? If needed, to what extent? This question quantifies the proportion of Agentic-PRs merged without any modifications and, if revisions are needed, the extent of such changes. Results show a similar proportion of Agentic-PR ($54.947\,37\text{\,}$%) and Human-PR ($58.527\,13\text{\,}$%) are accepted without revisions, and when revisions are required, the effort in terms of commits, lines of code, and files modified is not statistically different between the two groups.
@@ 2. Motivating Example
Claude Code is an AI-powered assistant designed to support a broad spectrum of software engineering tasks through natural language interaction. A key innovation of Claude Code is its agentic tool use [^2] https://www.anthropic.com/news/claude-3-7-sonnet and its integration with the Model Context Protocol (MCP), [^3] https://docs.anthropic.com/en/docs/claude-code/tutorials#set-up-model-context-protocol-mcp which significantly enhances its capabilities by connecting it with diverse data sources and tools, such as filesystems, command-line interfaces, and cloud services. As of May 2025, Claude Code and agents powered by Claude 4 Sonnet have outperformed several competing approaches on SWE-bench [^4] https://www.swebench.com (above 70% resolution rate), a benchmark for evaluating AI performance on real-world software issues on GitHub (Jimenez et al., 2024).
@@
Fig. 1 illustrates a command-line interaction in which Claude Code performs a refactoring task. Developers can instruct Claude Code to refactor the codebase to enhance readability and reduce cyclomatic complexity. In response, Claude analyzes the files, applies appropriate transformations, such as extracting helper methods, and updates the corresponding source file. When integrated with version control systems, Claude Code can then generate a pull request (i.e., an Agentic-PR) that packages the modifications for review. The PR description is automatically synthesized, documenting the changes made and indicating its provenance with the phrase “Generated with Claude Code”. Fig. 6 presents an example of an Agentic-PR on GitHub automatically created by Claude Code. This PR includes both the code modifications and a structured description summarizing the intent and scope of the change. Such automation lowers the effort required for developers to prepare contributions, but also raises questions about the quality and trustworthiness of these agent-generated artifacts.
Importantly, the quality of Agentic-PRs can vary. In some cases, they are immediately useful and merged with little modification. In other cases, developers express dissatisfaction. For instance, during the review process for an Agentic-PR, [^8] https://github.com/shopwareLabs/phpstan-shopware/pull/16 a reviewer identified an issue in the generated code, and the author noted that the proposed change can not be used at all. In the end, the PR is closed without merging. Even when code generated by agentic coding approaches is ultimately accepted, it often requires substantial revision by developers. For example, the bottom of Fig. 6 shows how developers made several follow-up commits to improve the code originally produced by Claude Code.
@@ 3. Data Collection
Additionally, to characterize PRs related to agent-assisted coding, we constructed a comparison set of human-created PRs (hereafter referred to as Human-PRs). Specifically, we selected PRs created during a similar time frame and by the same author, originating from the same project repositories. To ensure a balanced comparison, we randomly sampled Human-PRs to match the number of Agentic-PRs and their authors per repository. However, due to the limited availability of matching PRs, we extended the sampling window to one year for Human-PRs. We excluded 176 Agentic-PRs from the agent-assisted coding dataset because we could not obtain the same number of Human-PRs created during a similar time frame and by the same author. In the end, both datasets have 567 PRs across 157 repositories. Importantly, we ensured that the Agentic-PR and Human-PR datasets were mutually exclusive, with no overlapping PRs.
@@ 4.1. $RQ_{1}$: How do Agentic-PRs differ from Human PRs in terms of change size and purpose?
Approach. To investigate the purpose behind Agentic-PRs (APRs) and Human-PRs (HPRs), we sampled from a population of 475 merged APRs and 516 merged HPRs, and performed a manual classification on 213 APRs and 221 HPRs to understand the purpose of these PRs. We calculated the necessary sample size for each of these respective populations to satisfy a 95% confidence level with a 5% margin of error. Each PR was categorized using the two-dimensional framework proposed by Zeng et al. (Zeng et al., 2025), which distinguishes between purpose types (fix, feat, refactor, style, perf) and object types (docs, test, ci, build, chore). Although this framework was originally developed for commit-level classification, we apply it to PR-level analysis as PR descriptions typically summarize the collective purpose of commits. Therefore, we did not apply their hierarchical rule where purpose takes precedence over object when overlaps occur, and multiple labels can be applied to each PR. The initial classification, conducted independently by the first and third authors, achieved a label-level agreement of 75.8%. Specifically, disagreements occurred on 172 out of 675 labels for Agentic-PRs and 113 out of 501 labels for Human-PRs. Since our classification was a multi-label task, we report this agreement rate instead of Cohen’s Kappa (Landis and Koch, 1977). To resolve disagreements, the second author reviewed all conflicting cases and proposed final labels for each disputed case. The three authors then discussed until a unanimous consensus was reached on all annotations. The first three authors have 7, 15, and 13 years of programming experience, respectively, ensuring sufficient domain expertise for accurate classification.
@@
Results. Bug fixing is the most common category for both Agentic-PRs (31.031.0%) and Human-PRs (30.830.8%). Feature development is the second most common for both, accounting for 26.826.8% of Agentic-PRs and 27.627.6% of Human-PRs. As shown in Table 1, 66 Agentic-PRs and 68 Human-PRs were dedicated to bug fixing, making it the largest share for both. For example, an Agentic-PR  [^10] https://github.com/mattermost/mattermost/pull/30611 corrected previously ignored HTTP request handling errors and improved linter coverage, demonstrating how AI can accelerate debugging. In addition, Agentic-PRs are also frequently used for developing features (26.826.8%), including some domain-specific tasks. For example, an Agentic-PR implemented functionality to store both original audio files and their corresponding WAV transcription versions, enhancing playback and transcription quality. [^11] https://github.com/rishikanthc/Scriberr/pull/71
@@ 4.2. $RQ_{2}$: To what extent are Agentic-PRs rejected and why?
To categorize the reasons for PR rejections, we adopted the classification framework proposed by Pantiuchina et al. (Pantiuchina et al., 2022), which identifies rejection patterns in refactoring PRs through empirical analysis of code review comments. Although their work focused on refactoring, the framework applies to our broader context, as many rejection reasons are independent of the change type. The framework systematically organizes rejection reasons into process-related and refactoring-specific categories. We use five subcategories: (i) three from process-related subcategories called Are inactive (author/community), Are too large, and Are obsolete ; and (ii) two from refactoring-specific subcategories called Do not add value and Contain choices of non-optimal design solutions. The classification process by two independent authors achieved 68.5% agreement between the two primary annotators, with a Cohen’s K coefficient of 0.4865, indicating moderate agreement (Landis and Koch, 1977).
The selection of these categories was motivated by two factors: First, preliminary analysis of our dataset revealed that these rejection patterns occurred with high frequency. Second, employing an empirically-validated classification enhances the reliability of our findings and facilitates comparison with existing literature. However, during our analysis, we identified rejection reasons that could not be adequately classified within the framework from Pantiuchina et al. (Pantiuchina et al., 2022). For these cases, we inductively derived new categories through open coding of rejection comments, thereby extending the original classification to comprehensively capture the rejection patterns specific to our context.
@@
The most common reasons for Agentic-PR rejections stem from project evolution and PR complexity, not just code quality. Table 3 shows that alternative solutions (12.112.1%), obsolescence (3.33.3%), and oversized PRs (3.33.3%) frequently result in Agentic-PR rejections, reflecting project evolution and maintainability concerns. For example, one Agentic-PR  [^21] https://github.com/mattermost/mattermost/pull/30593 was closed after the team resolved the issue using a different solution, stating, “We might come back to this, but for now solving the underlying question in a different way.” Another PR [^22] https://github.com/OpenAdaptAI/OpenAdapt/pull/946 became obsolete after a newer contribution addressed the same functionality. Similarly, a large PR [^23] https://github.com/solvespace/solvespace/pull/1553 was closed with the comment, “Closing in favor of smaller, more focused PRs to make reviews more manageable,” emphasizing the difficulty of integrating oversized contributions into collaborative review processes.
@@
Table 3. Reasons for Agent PR rejection
@@
Technical shortcomings, such as issues with code implementation quality rather than strategic choices, account for 4.4% of Agentic-PR rejections, with 2.22.2% rejected for non-optimal design solutions, 1.11.1% for overly complex implementations, and 1.11.1% for introducing bugs or breaking functionality. Table 3 shows that 4.4% of the rejected Agentic-PRs are due to technical reasons where the implementation approach or code quality was problematic. For example, one PR [^26] https://github.com/zenml-io/zenml/pull/3375 was closed for bypassing project-specific serialization mechanisms, violating architectural principles. Another PR [^27] https://github.com/pytorch/test-infra/pull/6410 was rejected after reviewers found that existing themes could achieve the same functionality with reduced complexity. Furthermore, a PR [^28] https://github.com/pytorch/test-infra/pull/6400 closed with the comment “this is badly extracted and has some hallucinations, new CSS changes, and more importantly breaks the UI,” demonstrating functional failures. These examples highlight persistent limitations in AI’s ability to generate quality code that maintains system integrity.
Strategic misalignment, such as issues with what to build rather than how to build it, accounts for 2.2% of Agentic-PR rejections, with 1.11.1% closed for not adding value and 1.11.1% for not aligning with community interests. Table 3 indicates these PRs targeted wrong problems or proposed unnecessary changes, regardless of implementation quality. For instance, one PR [^29] https://github.com/mattermost/mattermost/pull/30890 was closed after maintainers determined the changes, while technically correct, failed to address the actual performance bottlenecks that mattered. Another PR [^30] https://github.com/MaiM-with-u/MaiBot/pull/195 was rejected after reviewers referenced the project’s policy document, determining the proposed large-scale changes did not provide sufficient benefit to justify their inclusion. These cases demonstrate challenges in AI’s understanding of project priorities and identifying which problems actually need solving.
Contributor inaction accounts for 2.22.2% of Agentic-PR rejections, with PRs abandoned when contributors stop responding to the review process. Table 3 indicates that 2 rejected APRs were closed due to general inactivity. For instance, a PR [^31] https://github.com/osmosis-labs/osmosis/pull/9029 was marked stale and automatically closed after the contributor failed to respond to implementation concerns. Another case [^32] https://github.com/ApeWorX/ape/pull/2532 involved a reviewer commenting “Waiting on this to be resolved” with a link to specific code review feedback, but the contributor never addressed the issue, resulting in automated closure by a bot. These examples illustrate how Agentic-PRs, when not actively managed by human contributors, face higher rejection risks due to minimal post-submission engagement.
@@ 4.3. $RQ_{3}$: What proportion of Agentic-PRs are accepted without revisions? If needed, to what extent?
When revisions are required, there is no statistically significant difference in revision frequency or cost. Analysis of revision patterns shows that both APRs and HPRs typically include a median of 2 revised commits prior to merge (see Fig. 5). The Mann-Whitney U tests show no statistically significant differences between Agentic-PRs and Human-PRs in terms of revision frequency. Although the median number of files, lines added, and lines deleted in revised commits can differ between the two groups (see Fig. 6), none of these differences are statistically significant at $\alpha=0.05$. These results suggest that the revision cost for Agentic-PR is not significantly different from that for Human-PR. Given this equivalence, leveraging AI to automate the initial stage of PR creation could effectively reduce developer burden and improve productivity.
@@
Approach. We examined all 475 merged Agentic-PRs and identified 214 Agentic-PRs (45.145.1%) that had modifications between initial submission and merge. We performed manual classification on these 214 Agentic-PRs to understand what types of changes were made during the review process. We manually inspect all commits added after the initial submission and classify them using the same two-dimensional framework from Zeng et al. (Zeng et al., 2025) described in Section 4.1. For each PR, we aggregated revision types at the PR level to avoid over-representing PRs with numerous small commits. For example, if a PR included three bug fixes and two stylistic changes across separate commits, it was counted as one instance of each change type. This manual classification was performed by the second and third authors. As in Section 4.1, we treated this as a multi-label task and measured agreement at the label level. The overall agreement rate was 64.9% (541 out of 844 labels). Disagreements were resolved by the second author following the same process described in Section 4.1. In addition, to understand the extent of continued AI assistance during the revision process, we analyzed co-authorship patterns in revision commits. We extracted commit metadata to identify commits containing “Co-Authored-By: Claude” attribution.
@@
New features are added in 14.614.6% of Agentic-PRs by human reviewers. Table 4 indicates that 33 out of 226 modified Agent PRs (14.614.6%) received feature additions during the review process. For example, one revision [^42] https://github.com/github/github-mcp-server/pull/118/commits/0527bc552591df807ee7be0b3a79f428d2f0ed2c expanded the PR review API functionality that the agent had initially implemented with basic capabilities. The revision added support for multi-line comments, including new parameters for line positioning, validation logic for parameter combinations, and comprehensive test coverage. This enhancement demonstrates that agent-generated implementations often provide core functionality but miss advanced features or edge cases that developers identify during review, requiring subsequent additions to meet complete user requirements.
Build configuration adjustments occur in 13.313.3% of reviewed Agent PRs. Table 4 shows that 30 out of 226 modified Agentic-PRs received build configuration updates during review. One revision [^43] https://github.com/coder/coder/pull/17035/commits/3c3aa219b895a981dd13a1b530603aa257f5549e addressed compatibility issues between the linting tool and the newer Go version. The agent’s initial implementation did not account for version incompatibilities, causing build failures with Go 1.24.1. The revision updated the linter version, configured it to run with a reduced set of checks, and modified the build process to continue despite linting failures. This temporary solution allowed the build pipeline to function while a comprehensive fix was planned for a subsequent update. The revision illustrates that agent-generated build configurations may overlook toolchain compatibility requirements, requiring manual intervention to maintain functional build processes across version updates.

### v2

@@ Abstract.
submissionid: Under minor revision at ACM TOSEMcopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Integrated and visual development environmentsccs: Software and its engineering Automatic programmingccs: Software and its engineering Software evolutionccs: Software and its engineering Maintaining software
This manuscript is currently under minor revision at ACM Transactions on Software Engineering and Methodology (TOSEM).
@@ 1. Introduction
Prior research has examined the impact of integrating LLM-based tools into software development workflows. For instance, Tufano et al. (DBLP:conf/msr/TufanoMPDPB24) manually analyzed 527 commits, 327 pull requests (PRs), and 647 issues, revealing that developers employ ChatGPT for a wide range of software engineering tasks, spanning 45 distinct categories, including automating the creation or enhancement of features. Similarly, Watanabe et al. (DBLP:conf/ease/WatanabeK0HYI24) studied the use of ChatGPT in PR reviews and the corresponding developer responses. They found that 30.7% of AI-generated suggestions during code reviews were met with skepticism or disagreement, and most of them were about the generated code.
@@
  • $RQ_{3}$: What proportion of Agentic-PRs are accepted without revisions? If needed, to what extent? This question quantifies the proportion of Agentic-PRs merged without any modifications and, if revisions are needed, the extent of such changes. Results show a similar proportion of Agentic-PRs ($54.947\,37\text{\,}$%) and Human-PRs ($58.527\,13\text{\,}$%) are accepted without revisions, and when revisions are required, the effort in terms of commits, lines of code, and files modified is not statistically different between the two groups.
@@ 2. Motivating Example
Claude Code is an AI-powered assistant designed to support a broad spectrum of software engineering tasks through natural language interaction. A key innovation of Claude Code is its agentic tool use [^2] https://www.anthropic.com/news/claude-3-7-sonnet and its integration with the Model Context Protocol (MCP), [^3] https://docs.anthropic.com/en/docs/claude-code/tutorials#set-up-model-context-protocol-mcp which significantly enhances its capabilities by connecting it with diverse data sources and tools, such as filesystems, command-line interfaces, and cloud services. As of May 2025, Claude Code and agents powered by Claude 4 Sonnet have outperformed several competing approaches on SWE-bench [^4] https://www.swebench.com (above 70% resolution rate), a benchmark for evaluating AI performance on real-world software issues on GitHub (DBLP:conf/iclr/JimenezYWYPPN24).
@@
Fig. 1 illustrates a command-line interaction in which Claude Code performs a refactoring task. Developers can instruct Claude Code to refactor the codebase to enhance readability and reduce cyclomatic complexity. In response, Claude analyzes the files, applies appropriate transformations, such as extracting helper methods, and updates the corresponding source file. When integrated with version control systems, Claude Code can then generate a pull request (i.e., an Agentic-PRs) that packages the modifications for review. The PR description is automatically synthesized, documenting the changes made and indicating its provenance with the phrase “Generated with Claude Code”. Fig. 6 presents an example of an Agentic-PRs on GitHub automatically created by Claude Code. This PR includes both the code modifications and a structured description summarizing the intent and scope of the change. Such automation lowers the effort required for developers to prepare contributions, but also raises questions about the quality and trustworthiness of these agent-generated artifacts.
Importantly, the quality of Agentic-PRs can vary. In some cases, they are immediately useful and merged with little modification. In other cases, developers express dissatisfaction. For instance, during the review process for an Agentic-PRs, [^8] https://github.com/shopwareLabs/phpstan-shopware/pull/16 a reviewer identified an issue in the generated code, and the author noted that the proposed change can not be used at all. In the end, the PR is closed without merging. Even when code generated by agentic coding approaches is ultimately accepted, it often requires substantial revision by developers. For example, the bottom of Fig. 6 shows how developers made several follow-up commits to improve the code originally produced by Claude Code.
@@ 3. Data Collection
Additionally, to characterize PRs related to agent-assisted coding, we constructed a comparison set of human-created PRs (hereafter referred to as Human-PRs). Specifically, we selected PRs created during a similar time frame and by the same author, originating from the same project repositories. To ensure a balanced comparison, we randomly sampled Human-PRs to match the number of Agentic-PRs and their authors per repository. However, due to the limited availability of matching PRs, we extended the sampling window backward for Human-PRs until the sample size matched that of Agentic-PRs. This extension resulted in a collection period spanning approximately 10 months (back to April 30, 2024). We excluded 176 Agentic-PRs from the agent-assisted coding dataset because we could not obtain the same number of Human-PRs created during a similar time frame and by the same author. In the end, both datasets have 567 PRs across 157 repositories. Importantly, we ensured that the Agentic-PR and Human-PR datasets were mutually exclusive, with no overlapping PRs.
@@ 4.1. $RQ_{1}$: How do Agentic-PRs differ from Human PRs in terms of change size and purpose?
Approach. To investigate the purpose behind Agentic-PRs (APRs) and Human-PRs (HPRs), we sampled from a population of 475 merged APRs and 516 merged HPRs, and performed a manual classification on 213 APRs and 221 HPRs to understand the purpose of these PRs. We calculated the necessary sample size for each of these respective populations to satisfy a 95% confidence level with a 5% margin of error. Each PR was categorized using the two-dimensional framework proposed by Zeng et al. (DBLP:conf/icse/ZengZQL25), which distinguishes between purpose types (fix, feat, refactor, style, perf) and object types (docs, test, ci, build, chore). Although this framework was originally developed for commit-level classification, we apply it to PR-level analysis as PR descriptions typically summarize the collective purpose of commits. Therefore, we did not apply their hierarchical rule where purpose takes precedence over object when overlaps occur, and multiple labels can be applied to each PR. The initial classification, conducted independently by the first and third authors, achieved a label-level agreement of 75.8%. Specifically, disagreements occurred on 172 out of 675 labels for Agentic-PRs and 113 out of 501 labels for Human-PRs. Since our classification was a multi-label task, we report this agreement rate instead of Cohen’s Kappa (kappa). To resolve disagreements, the second author reviewed all conflicting cases and proposed final labels for each disputed case. The three authors then discussed until a unanimous consensus was reached on all annotations. The first three authors have 7, 15, and 13 years of programming experience, respectively, ensuring sufficient domain expertise for accurate classification.
@@
Results. Bug fixing is the most common category for both Agentic-PRs (31.031.0%) and Human-PRs (30.830.8%). Feature development is the second most common for both, accounting for 26.826.8% of Agentic-PRs and 27.627.6% of Human-PRs. As shown in Table 1, 66 Agentic-PRs and 68 Human-PRs were dedicated to bug fixing, making it the largest share for both. For example, an Agentic-PRs  [^10] https://github.com/mattermost/mattermost/pull/30611 corrected previously ignored HTTP request handling errors and improved linter coverage, demonstrating how AI can accelerate debugging. In addition, Agentic-PRs are also frequently used for developing features (26.826.8%), including some domain-specific tasks. For example, an Agentic-PRs implemented functionality to store both original audio files and their corresponding WAV transcription versions, enhancing playback and transcription quality. [^11] https://github.com/rishikanthc/Scriberr/pull/71
@@ 4.2. $RQ_{2}$: To what extent are Agentic-PRs rejected and why?
To categorize the reasons for PR rejections, we adopted the classification framework proposed by Pantiuchina et al. (DBLP:journals/tosem/PantiuchinaLZPL22), which identifies rejection patterns in refactoring PRs through empirical analysis of code review comments. Although their work focused on refactoring, the framework applies to our broader context, as many rejection reasons are independent of the change type. The framework systematically organizes rejection reasons into process-related and refactoring-specific categories. We use five subcategories: (i) three from process-related subcategories called Are inactive (author/community), Are too large, and Are obsolete ; and (ii) two from refactoring-specific subcategories called Do not add value and Contain choices of non-optimal design solutions. The classification process by two independent authors achieved 68.5% agreement between the two primary annotators, with a Cohen’s K coefficient of 0.4865, indicating moderate agreement (kappa).
The selection of these categories was motivated by two factors: First, preliminary analysis of our dataset revealed that these rejection patterns occurred with high frequency. Second, employing an empirically-validated classification enhances the reliability of our findings and facilitates comparison with existing literature. However, during our analysis, we identified rejection reasons that could not be adequately classified within the framework from Pantiuchina et al. (DBLP:journals/tosem/PantiuchinaLZPL22). For these cases, we inductively derived new categories through open coding of rejection comments, thereby extending the original classification to comprehensively capture the rejection patterns specific to our context.
@@
The most common reasons for Agentic-PRs rejections stem from project evolution and PR complexity, not just code quality. Table 3 shows that alternative solutions (12.112.1%), obsolescence (3.33.3%), and oversized PRs (3.33.3%) frequently result in Agentic-PRs rejections, reflecting project evolution and maintainability concerns. For example, one Agentic-PRs  [^21] https://github.com/mattermost/mattermost/pull/30593 was closed after the team resolved the issue using a different solution, stating, “We might come back to this, but for now solving the underlying question in a different way.” Another PR [^22] https://github.com/OpenAdaptAI/OpenAdapt/pull/946 became obsolete after a newer contribution addressed the same functionality. Similarly, a large PR [^23] https://github.com/solvespace/solvespace/pull/1553 was closed with the comment, “Closing in favor of smaller, more focused PRs to make reviews more manageable,” emphasizing the difficulty of integrating oversized contributions into collaborative review processes.
@@
Table 3. Reasons for Agentic-PRs rejection
@@
Technical shortcomings, such as issues with code implementation quality rather than strategic choices, account for 4.4% of Agentic-PRs rejections, with 2.22.2% rejected for non-optimal design solutions, 1.11.1% for overly complex implementations, and 1.11.1% for introducing bugs or breaking functionality. Table 3 shows that 4.4% of the rejected Agentic-PRs are due to technical reasons where the implementation approach or code quality was problematic. For example, one PR [^26] https://github.com/zenml-io/zenml/pull/3375 was closed for bypassing project-specific serialization mechanisms, violating architectural principles. Another PR [^27] https://github.com/pytorch/test-infra/pull/6410 was rejected after reviewers found that existing themes could achieve the same functionality with reduced complexity. Furthermore, a PR [^28] https://github.com/pytorch/test-infra/pull/6400 closed with the comment “this is badly extracted and has some hallucinations, new CSS changes, and more importantly breaks the UI,” demonstrating functional failures. These examples highlight persistent limitations in AI’s ability to generate quality code that maintains system integrity.
Strategic misalignment, such as issues with what to build rather than how to build it, accounts for 2.2% of Agentic-PRs rejections, with 1.11.1% closed for not adding value and 1.11.1% for not aligning with community interests. Table 3 indicates these PRs targeted wrong problems or proposed unnecessary changes, regardless of implementation quality. For instance, one PR [^29] https://github.com/mattermost/mattermost/pull/30890 was closed after maintainers determined the changes, while technically correct, failed to address the actual performance bottlenecks that mattered. Another PR [^30] https://github.com/MaiM-with-u/MaiBot/pull/195 was rejected after reviewers referenced the project’s policy document, determining the proposed large-scale changes did not provide sufficient benefit to justify their inclusion. These cases demonstrate challenges in AI’s understanding of project priorities and identifying which problems actually need solving.
Contributor inaction accounts for 2.22.2% of Agentic-PRs rejections, with PRs abandoned when contributors stop responding to the review process. Table 3 indicates that 2 rejected APRs were closed due to general inactivity. For instance, a PR [^31] https://github.com/osmosis-labs/osmosis/pull/9029 was marked stale and automatically closed after the contributor failed to respond to implementation concerns. Another case [^32] https://github.com/ApeWorX/ape/pull/2532 involved a reviewer commenting “Waiting on this to be resolved” with a link to specific code review feedback, but the contributor never addressed the issue, resulting in automated closure by a bot. These examples illustrate how Agentic-PRs, when not actively managed by human contributors, face higher rejection risks due to minimal post-submission engagement.
@@ 4.3. $RQ_{3}$: What proportion of Agentic-PRs are accepted without revisions? If needed, to what extent?
When revisions are required, there is no statistically significant difference in revision frequency or cost. Analysis of revision patterns shows that both APRs and HPRs typically include a median of 2 revised commits prior to merge (see Fig. 5). The Mann-Whitney U tests show no statistically significant differences between Agentic-PRs and Human-PRs in terms of revision frequency. Although the median number of files, lines added, and lines deleted in revised commits can differ between the two groups (see Fig. 6), none of these differences are statistically significant at $\alpha=0.05$. These results suggest that the revision cost for Agentic-PRs is not significantly different from that for Human-PR. Given this equivalence, leveraging AI to automate the initial stage of PR creation could effectively reduce developer burden and improve productivity.
@@
Approach. We examined all 475 merged Agentic-PRs and identified 214 Agentic-PRs (45.145.1%) that had modifications between initial submission and merge. We performed manual classification on these 214 Agentic-PRs to understand what types of changes were made during the review process. We manually inspect all commits added after the initial submission and classify them using the same two-dimensional framework from Zeng et al. (DBLP:conf/icse/ZengZQL25) described in Section 4.1. For each PR, we aggregated revision types at the PR level to avoid over-representing PRs with numerous small commits. For example, if a PR included three bug fixes and two stylistic changes across separate commits, it was counted as one instance of each change type. This manual classification was performed by the second and third authors. As in Section 4.1, we treated this as a multi-label task and measured agreement at the label level. The overall agreement rate was 64.9% (541 out of 844 labels). Disagreements were resolved by the second author following the same process described in Section 4.1. In addition, to understand the extent of continued AI assistance during the revision process, we analyzed co-authorship patterns in revision commits. We extracted commit metadata to identify commits containing “Co-Authored-By: Claude” attribution.
@@
New features are added in 14.614.6% of Agentic-PRs by human reviewers. Table 4 indicates that 33 out of 226 modified Agentic-PRs (14.614.6%) received feature additions during the review process. For example, one revision [^42] https://github.com/github/github-mcp-server/pull/118/commits/0527bc552591df807ee7be0b3a79f428d2f0ed2c expanded the PR review API functionality that the agent had initially implemented with basic capabilities. The revision added support for multi-line comments, including new parameters for line positioning, validation logic for parameter combinations, and comprehensive test coverage. This enhancement demonstrates that agent-generated implementations often provide core functionality but miss advanced features or edge cases that developers identify during review, requiring subsequent additions to meet complete user requirements.
Build configuration adjustments occur in 13.313.3% of reviewed Agentic-PRs. Table 4 shows that 30 out of 226 modified Agentic-PRs received build configuration updates during review. One revision [^43] https://github.com/coder/coder/pull/17035/commits/3c3aa219b895a981dd13a1b530603aa257f5549e addressed compatibility issues between the linting tool and the newer Go version. The agent’s initial implementation did not account for version incompatibilities, causing build failures with Go 1.24.1. The revision updated the linter version, configured it to run with a reduced set of checks, and modified the build process to continue despite linting failures. This temporary solution allowed the build pipeline to function while a comprehensive fix was planned for a subsequent update. The revision illustrates that agent-generated build configurations may overlook toolchain compatibility requirements, requiring manual intervention to maintain functional build processes across version updates.